$SPAD/src/input e02agf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1996.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
m:=5
kplus1:=5
nrows:=6
xmin:=0.0
xmax:=4.0
x:Matrix SF:=
 [[0.5 ,1.0 ,2.0 ,2.5 ,3.0 ]]
y:Matrix SF:=
 [[0.03 ,-0.75 ,-1.0 ,-0.1 ,1.75 ]]
w:Matrix SF:=
 [[1.0 ,1.0 ,1.0 ,1.0 ,1.0 ]]
mf:=2
xf:Matrix SF:=
 [[0.0 ,4.0 ]]
yf:Matrix SF:=
 [[1.0 ,-2.0 ,9.0 ,0.0 ,0.0 ,0.0 ,_
   0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,_
   0.0 ,0.0 ,0.0 ]]
lyf:=15
ip:Matrix Integer:=
 [[1 ,0 ]]
lwrk:=200
liwrk:=12
result:=e02agf(m,kplus1,nrows,xmin,xmax,x,y,w,mf,xf,yf,lyf,ip,lwrk,liwrk, -1)

———-

References^{[1] nothing}

4